A body falling from a vertical height of 10 m pierces through a distance of 1 m in sand. it faces an average retardation equal to (g= acceleration due to gravity):
1)g2)9g
3)100g
4)1000g

Grade:11

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hi
v^2=2gs=>200
again
v^2=2as=>a=200/2=100m/s^2
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
KAPIL MANDAL
9 years ago
Thank you sir for this answer. Thank you very much.